# LLMs Process Lists With General Filter Heads

### v1

@@ 2.2 Filter Heads
We observe that, for a range of filtering tasks, specific attention heads in the middle layers of Llama-70B consistently focus their attention on the items satisfying the given predicate, $\psi$. See Figure˜1 (more in Appendix˜M) where we show the attention distribution for these filter heads from the last token position. From Equation˜1, we know that this selective attention pattern emerges from the interaction between the query state at the last token ($q_{-1}$) and the key states from all preceding tokens ($K_{\leq t}=\{k_{1},k_{2},\ldots,k_{t}\}$). We employ activation patching (Meng et al., 2022; Zhang & Nanda, 2024) to understand the distinct causal roles of these states.
@@ 2.3 Locating Filter Heads
Activation Patching with DCM.
@@
Notably, while finding the heads we do not care if the heads exhibit the attention behavior illustrated in Figure˜1. But, we notice that the aggregated attention pattern of the identified heads consistently aligns with the selective attention pattern for filtering (see Appendix˜M for some examples).
@@ 3 Experiments
We study autoregressive transformer LMs in our experiments. Unless stated otherwise, all the reported results are for Llama-70B (Touvron et al., 2023). We include additional results for Gemma-27B (Team et al., 2024) in Appendix˜K.
@@ 3.2 Portability/Generalizability Across Filter-Reduce Operations
Figure 3: Generalization across different tasks. (a) shows whether the heads identified with one task (rows) maintain causal influence in another task (columns). (b) shows how portable the predicate representation is across tasks. The predicate rep $q_{\text{src}}$ is cached from one source task example (e.g., find the fruit in SelectOne task) and was patched to an example from another destination task (e.g., count the vehicles in Counting task). The heatmap shows causality scores, i.e. whether the LM correctly performs the destination task with the transferred predicate (e.g., count the fruits). For both (a) and (b) the values in the diagonal grid show within task scores.
@@ 3.3 Necessity of Filter Heads
Concept $0.080$ +5.23+5.23
Induction $0.00$ 3.23-3.23
@@
To determine whether filter heads represent a novel discovery or merely overlap with existing attention head categories previously documented in the literature, we compared their functionality against such head categories. Specifically, we measure the causality of Function Vector heads (Todd et al., 2024), Concept heads (Feucht et al., 2025), and Induction heads (Olsson et al., 2022). As shown in Table 4, none of these previously identified head types exhibit the distinctive functional role of filter heads, confirming that filter heads are a unique and previously unrecognized component of transformer LMs.
@@ 4 Key States Carry Item Semantics for Filtering
This result confirms our mechanistic hypothesis: filter heads implement filtering through a key-query interaction where queries encode what to look for (the predicate) and keys bring what is there (item properties) from the corresponding item latents ($h^{\ell-1}$).
@@ 5 What happens if the question comes before the options?
In Section˜3.1(c) we see that if we simply reverse the order of the question and the collection to ask the question before presenting the items, the causality scores drop to almost zero. Our investigations reveal that this seemingly innocent ordering change fundamentally alters the computational strategy available to the LM. When the question comes first, the transformer can perform eager evaluation: as each item is processed, the model can immediately evaluate whether it satisfies the predicate and store this information as an is_match flag in the item’s latents. And, at the final token, rather than performing the predicate matching operation via filter heads, the LM can simply retrieve items based on pre-computed flags. If this hypothesized flagging mechanism is true, then manipulating this flag should result in predictable outcomes in the LM’s behavior. We find evidence for this alternative mechanism through a series of carefully designed activation patching experiments. We illustrate the core experiment setup in Figure˜5, while we leave the detailed analysis to Appendix˜B.
@@
However, filter heads remain partially active even in question-first examples. Section˜3.1(c) also shows that caching $q_{\mathrm{src}}$ from question-first prompts and patching to question-after prompts gives non-trivial causality scores, though lower than our original setup. This suggests filter heads can still partially encode the predicate information in question-first settings, but this filter head based mechanism competes with, and is typically overshadowed by, the flag-based mechanism.
The co-existence of these two filtering strategies: on-demand filtering through filter heads and precomputing the flags and storing them in the item latents, echoes the lazy versus eager evaluation strategies in functional programming from Henderson & Morris Jr (1976). This also shows how transformer LMs can maintain multiple redundant pathways for the same operation (McGrath et al., 2023; Wang et al., 2023) and can dynamically select between them based on task demands and information availability.
@@ 6 Application: A light-weight Probe Without Training
Since filter heads encode predicates as query states that interact with key-projected item representations to perform filtering, we can repurpose this mechanism for classification. To detect whether a representation $h$ belongs to a particular concept class (e.g., animal, vehicle), we create filter prompts for each class: $p_{\text{cls}}=\mathbb{P}(\mathcal{C},\psi_{\mathrm{is\_cls}})$ and collect the query states $q_{\mathrm{cls}}$ from a filter head $\ell,j]$. Then we classify $h$ by finding the class whose $q_{\mathrm{cls}}$ has the maximum affinity.
@@
Where $W_{K}^{\ell j}$ is the key projection from the head $\ell,j]$. Figure 6 demonstrates that this approach achieves strong classification performance without any training, validating that filter heads learn generalizable concept representations that can be extracted and applied as probes. See Section˜M.3 where we illustrate how filter heads can be utilized to detect other concepts such as the presence of false information or certain sentiment in free-form text.
@@ 8 Discussion
In this work, we have identified and characterized filter heads — specialized attention heads that implement filtering operations in autoregressive transformer LMs. These heads encode the filtering criteria (predicates) as compact representations in their query states of specific tokens. This encoding can be extracted and then transported to another context to trigger the same operation. We also identify that, based on information availability, the LM can use an eager implementation of filtering by storing flags directly on the item latents. These dual and complimentary filtering implementations mirror the lazy vs eager evaluation from functional programming. This convergence between emergent neural mechanisms and human-designed programming primitives suggests that certain computational patterns arise naturally from task demands rather than architectural constraints. Cataloging such universal computational primitives and how they are realized may help us understand how AI systems perform complex reasoning.
@@
The code and dataset produced in this work are available at github.com/arnab-api/filter. We ran all experiments on workstations with either 80GB NVIDIA A100 GPUs or 48GB A6000 GPUs, using the HuggingFace Transformers library (Wolf et al., 2019) and PyTorch (Paszke et al., 2019). We used NNsight (Fiotto-Kaufman et al., 2025) for our patching experiments.
@@ Acknowledgements
This research has been supported by a grant from Open Philanthropy (DB, AS, NS), and the Israel Council for Higher Education (NS). We have also received compute credits from the U.S. NSF Advanced Cyberinfrastructure Coordination Ecosystem: Services & Support (NSF ACCESS), which supported some of our experiments.
@@ A.2.1 Item Presentation
Single Line ‘‘Options: House, Blender, Willow, Ambulance, Piano, Wrestling mat.
@@ Appendix B Dual Implementation of Filtering in LMs: Question Before vs After
Section˜3.1(c) shows that filter heads are minimally causal when patched from question-before to question-before prompt, even when both follow the same prompt template and item presentation style. To understand this better, we perform a multi-step causal mediation analysis.
@@ 3. O3:
To test this hypothesis, we ablate the is_match flag in an item by replacing all the residual stream latents of the tokens of that item with their corresponding latents cached for the same item in a neutral prompt (see Figure˜9 for details).
@@
When we perform this ablation for each of the items in the collection, we indeed see a significant drop in LM performance for the question-before setting, while the performance remains mostly unchanged for the question-after setting (see Table˜5).
This experiment supports our hypothesis that the question-before setting allows the LM to eagerly evaluate whether an item satisfies the predicate or not, store this intermediate result in the residual latents of the items as it processes the collection, and relies on that to make the final decision. We also notice that the question-after setting shows minimal sensitivity to this flag-ablation, which suggests that the processing of items do not rely on the context here: the LM populates the semantics (enrichment in Geva et al. (2023)) of each item in a context independent manner first, and then applies the predicate to perform filtering when the question is presenter after.
@@
Our most decisive evidence comes from swapping the is_match flag between items: if we swap the is_match flag stored in $c_{\text{pos}}$ that satisfies the predicate with $c_{\text{neg}}$ that does not satisfy the predicate, we should expect the LM to change its answer from $c_{\text{pos}}$ to $c_{\text{neg}}$ if it is relying on the is_match flag. To test this hypothesis we set up another activation patching experiment (illustrated in Figure˜10, which is a more elaborate version of Figure˜5 in the main text). We perform a 2 part intervention:
@@ 2. I2:
Figure 11: Effect of swapping the is_match flag between items. The figures on the left shows the effect of patching only the residual states (Figure˜7) and figures on the right are when we additionally swap the is_match with another item (Figure˜10). The pair of figures on the top shows both cases in the question-before format (a). We observe that $c_{\text{flag}}$ becomes of top scoring item when the patching is performed in early layers. But this swapping of flags has no effect in the question after case, pair of figures on the bottom (b).
@@
We do not claim that the LM only relies on the is_match flag in the question-before setting. The fact that we see a bump in the score of $c_{\mathrm{targ}}$ ( blue) in Figure˜8 and patching $q_{\mathrm{src}}$ from question-before to question-after prompts has non-trivial causality (see Section˜3.1c) indicates that the LM does carry the predicate information in middle layers even in the question-before setting, although not strongly as in the question-after setting.
This dual filtering implementation strategy — lazy evaluation via filter heads versus storing eager evaluation information with is_match flag — exemplifies a broader principle in neural computation: transformer LMs can maintain multiple pathways for core operations, dynamically selecting strategies based on what information is available. And the fact that filter heads still remain partially active even in the question-first setting shows that these mechanisms operate in parallel rather than mutual exclusion.
@@ Appendix C Different Approaches for Locating the Filter Heads
$\text{FilterScore}([\ell,j],\mathcal{C},\psi)$ $=\text{score}_{\ell j}(c\mid\psi(c))-\max_{\neg\psi(c)}\;\big(\text{score}_{\ell j}(c)\big)$ (6)
$\text{where,}\quad\text{score}_{\ell j}(c)$ $=\sum_{t\in c}\text{Attn}_{[\ell,j]}(q_{-1},K)_{t}$
@@
Figure 12: Collective causality of 79 filter heads identified with FilterScore, CMA, and CMA with DCM. Evaluated on the same 512 examples from the SelectOne task
@@
Activation Patching (w/o DCM).
We can patch the attention heads individually and quantify their indirect effect at mediating the target property in the patched run. Todd et al. (2024) and Feucht et al. (2025) identified Function Vector heads and Concept heads with this approach. Specifically, for each head $\ell,j]$ we patch $q_{\mathrm{src}}$ from the source run to the destination run with the method detailed in Section˜2 and check its effect on boosting the score (logit) of the target item, $c_{\text{targ}}$. The indirect effect is measured as $\text{logit}[\leftarrow q_{\text{src}}]\big(c_{\text{targ}}\big)-\text{logit}\big(c_{\text{targ}}\big)$.
In Figure˜12 we compare the causality of heads identified with the 3 approaches on the SelectOne task.
Figure 13: Location of filter heads in Llama-70B. (a) shows the individual FilterScore for each head: how much they attend to the correct option over others. (b) shows the indirect effect: how much patching $q_{\mathrm{src}}$ from a single head promote the predicate target. The filter heads identified with Section˜2.3 are marked with black borders.

### v3

@@ 2.2 Filter Heads
We observe that, for a range of filtering tasks, specific attention heads in the middle layers of Llama-70B consistently focus their attention on the items satisfying the given predicate, $\psi$. See Figure˜1 (more in Appendix˜P) where we show the attention distribution for these filter heads from the last token position. From Equation˜1, we know that this selective attention pattern emerges from the interaction between the query state at the last token ($q_{-1}$) and the key states from all preceding tokens ($K_{\leq t}=\{k_{1},k_{2},\ldots,k_{t}\}$). We employ activation patching (Meng et al., 2022; Zhang & Nanda, 2024) to understand the distinct causal roles of these states.
@@ 2.3 Locating Filter Heads
Activation Patching with Learned Masks.
@@
Notably, while finding the heads we do not care if the heads exhibit the attention behavior illustrated in Figure˜1. But, we notice that the aggregated attention pattern of the identified heads consistently aligns with the selective attention pattern for filtering (see Appendix˜P for some examples).
@@ 3 Experiments
We study autoregressive transformer LMs in our experiments. Unless stated otherwise, all the reported results are for Llama-70B (Touvron et al., 2023). We include additional results for Gemma-27B (Team et al., 2024) in Appendix˜L.
@@ 3.2 Portability/Generalizability Across Filter-Reduce Operations
Figure 3: Generalization across different tasks. (a) shows whether the heads identified with one task (rows) maintain causal influence in another task (columns). (b) shows how portable the predicate representation is across tasks. The predicate rep $q_{\text{src}}$ is cached from one source task example (e.g., find the fruit in SelectOne task) and is patched to an example from another destination task (e.g., count the vehicles in Counting task). The heatmap shows causality scores, i.e. whether the LM correctly performs the destination task with the transferred predicate (e.g., count the fruits). For both (a) and (b) the values in the diagonal grid show within task scores.
@@ 3.3 Necessity of Filter Heads
Concept Induction $0.080$ +5.23+5.23
Token Induction $0.00$ 3.23-3.23
@@
To determine whether filter heads represent a novel discovery or merely overlap with existing attention head categories previously documented in the literature, we compared their functionality against such head categories with specific functional roles. Specifically, we measure the causality of Function Vector heads (Todd et al., 2024), Concept Induction heads (Feucht et al., 2025), and Token Induction heads (Olsson et al., 2022). As shown in Table 4, none of these previously identified head types exhibit the distinctive functional role of filter heads. We do not treat these heads as “baselines” in the traditional sense, as they are not intended for filtering tasks. We evaluate them to confirm that filter heads are a unique and previously unrecognized component in transformer LMs.
We also verify that filter heads encode abstract predicates rather than specific answer tokens: even when source prompts contain the predicate but no valid answer item in the collection, causality remains high at 0.80 (see Appendix E). Moreover, these predicate representations support composition through vector arithmetic: adding the query vectors of two predicates (e.g., is_fruit and is_vehicle) produces a vector that executes their disjunction, achieving a causality of 0.65 (see Appendix D). These results further confirm that filter heads encode predicates as compact geometric directions in query space.
@@ 4 Key States Carry Item Semantics for Filtering
These results confirm our mechanistic hypothesis: filter heads implement filtering through a localized key-query interaction where queries selectively encode what to look for (the predicate) while discarding the item content and other unrelated prompt context, and key states similarly bring what is in the list by extracting the semantic properties of the items from the corresponding latents ($h^{\ell-1}$), discarding the surrounding context. Notably, this behavior manifests in the attention pattern. Although previous works cautioned that attention patterns can be misleading (Jain & Wallace, 2019), in our predicate-matching setting we see a clear correspondence between how filter heads utilize abstract predicate encodings to attend to what was asked for and how they causally influence the LM behavior — doing what Vaswani et al. (2017) and the original designers of attention might have envisioned.
@@ 5 What happens if the question comes before the options?
In Section˜3.1(c) we see that if we simply reverse the order of the question and the collection to ask the question before presenting the items, the causality scores drop to near zero. Our investigations reveal that this seemingly innocent ordering change fundamentally alters the computational strategy available to the LM.
Hypothesis.
When the question comes first, the transformer can perform eager evaluation: as each item is processed, the model can immediately evaluate whether it satisfies the predicate and store this information as an is_match flag directly in the item’s latents. And, at the final token position, rather than performing the predicate matching operation via filter heads, the LM simply retrieves items based on these pre-computed flags.
If this hypothesized flagging mechanism is true, then manipulating this flag should result in predictable outcomes in the LM’s behavior. We find evidence for this alternative mechanism through a series of carefully designed causal mediation analysis. We illustrate the most decisive experiment setup in Figure˜5, while we leave the detailed analysis to Appendix˜B.
@@
Experiment Setup.
Figure˜5 illustrates our two-part causal intervention setup to test for the presence of the hypothesized flag information in the item latents. Similar to the setup described in Section˜2.2, we consider two prompts, a source prompt $p_{\mathrm{src}}=\mathbb{P}(\mathcal{C},\psi_{\mathrm{src}})$ and a destination prompt $p_{\mathrm{dest}}=\mathbb{P}(\mathcal{C},\psi_{\mathrm{dest}})$, both following either the question-before or question-after format. As in Section˜2.2, the predicates are different $(\psi_{\mathrm{src}}\neq\psi_{\mathrm{dest}})$, but they operate on the same collection of items in both the source and the destination prompt $(\mathcal{C}_{\mathrm{src}}=\mathcal{C}_{\mathrm{dest}}=\mathcal{C})$. Let $c_{\mathrm{src}}$ and $c_{\mathrm{dest}}$ be the items in $\mathcal{C}$ satisfying the predicates $\psi_{\mathrm{src}}$ and $\psi_{\mathrm{dest}}$ respectively.
Our goal is to engineer a situation where a third item $c_{\mathrm{flag}}\not\in\{c_{\mathrm{src}},c_{\mathrm{dest}}\}$ is the sole carrier of the is_match flag, and then check whether the LM is misled into selecting $c_{\mathrm{flag}}$. We achieve this with two simultaneous interventions:
  1. (a)
[…]
  2. (b)
[…]
Results.
Figure˜5 (right panel) shows the logit of different items as we vary the layer $\ell$ at which intervention (a) is applied, while intervention (b) is applied across all the layers throughout. In the question-before setting, the LM consistently selects $c_{\mathrm{flag}}$ in early-to-middle layers, which is precisely the behavior predicted by the flag-based hypothesis. The LM ignores both the source predicate ($\psi_{\mathrm{src}}$) and the destination predicate ($\psi_{\mathrm{dest}}$), and follows the planted flag instead. In contrast, replicating this experiment in the question-after setting shows that the LM is not strongly influenced by the flag, which is expected as the LM didn’t get an opportunity to precalculate the flag in the item latents when the predicate is presented after the items. See Appendix˜B for details on this.
Filter heads remain partially active.
Although the flag-based mechanism is dominant in the question-before setting, filter heads are not entirely inactive. Table 2(c) shows that caching $q_{\mathrm{src}}$ from question-before prompts and patching to question-after prompts yields non-trivial causality ($0.398$), indicating that filter heads still partially encode predicate information in question-first prompts. The two mechanisms thus operate in parallel rather than in mutual exclusion, with the flag-based pathway typically dominating when the question is presented before the collection.
@@ 6 Application: A light-weight Probe Without Training
Since filter heads encode predicates as query states that interact with key-projected item representations to perform filtering, we can repurpose this mechanism for classification. To detect whether a representation $h$ belongs to a particular concept class (e.g., animal, vehicle), we create filter prompts for each class: $p_{\mathrm{cls}}=\mathbb{P}(\mathcal{C},\psi_{\mathrm{is\_cls}})$ and collect the query states $q_{\mathrm{cls}}$ from a filter head $\ell,j]$.
Then we classify $h$ by finding the class whose $q_{\mathrm{cls}}$ has the maximum affinity.
@@
Where $W_{K}^{\ell j}$ is the key projection from the head $\ell,j]$. Figure 6 demonstrates that this approach achieves strong classification performance without any training, validating that filter heads learn generalizable concept representations that can be extracted and applied as probes.
Filter heads encode predicates in compact geometric directions that produce distinctive attention patterns via their interaction with the key states. We can use this insight to utilize filter heads to identify false information or detect sentiment in free-form text. See Section˜P.4 for some examples.
@@ 8 Discussion
In this work, we have identified and characterized filter heads — specialized attention heads that implement filtering operations in autoregressive transformer LMs. These heads encode the filtering criteria (predicates) as compact representations in their query states of specific tokens. This encoding can be extracted and then transported to another context to trigger the same operation. We also identify that, based on information availability, the LM can use an eager implementation of filtering by storing flags directly on the item latents. These dual and complimentary filtering implementations mirror the lazy vs eager evaluation from functional programming. This convergence between emergent neural mechanisms and human-designed programming primitives suggests that certain computational patterns arise naturally from task demands rather than architectural constraints.
@@
The code and dataset produced in this work are available at github.com/arnab-api/filter. We ran all experiments on workstations with either 80GB NVIDIA A100 GPUs or 48GB A6000 GPUs, using the HuggingFace Transformers library (Wolf et al., 2019) and PyTorch (Paszke et al., 2019). We used NNsight (Fiotto-Kaufman et al., 2025) for our intervention experiments.
@@ Acknowledgements
This research has been supported by a grant from Open Philanthropy (DB, AS, NS), and the Israel Council for Higher Education (NS). AS received compute credits from the U.S. NSF Advanced Cyberinfrastructure Coordination Ecosystem: Services & Support (NSF ACCESS), which supported some of our experiments.
@@ A.2.1 Item Presentation
Single Line ‘‘Options: House, Blender, Willow, Truck, Piano, Wrestling mat.
@@ A.2.3 From a different language
A.3 LM Baseline Performance on Tasks
Table 5: LLM baseline performance on filtering tasks. All the models were evaluated on the same draw of 1024 examples for each of the tasks. The LMs are sorted in descending order by size. The question was presented after the items.
Model (Huggingface ID)
SelectOne
SelectOne-MCQ
SelectFirst
SelectLast
Counting
CheckPresence
Llama-70B (meta-llama/Llama-3.3-70B-Instruct) $99.32\%$ $98.63\%$ $78.42\%$ $87.01\%$ $67.38\%$ $91.02\%$
Gemma-27B (google/gemma-2-27b-it) $99.02\%$ $97.75\%$ $75.29\%$ $88.96\%$ $64.16\%$ $96.29\%$
Gemma-9B (google/gemma-2-9b-it) $98.93\%$ $97.85\%$ $68.55\%$ $78.03\%$ $63.67\%$ $95.90\%$
Llama-8B (meta-llama/Llama-3.1-8B-Instruct) $98.44\%$ $97.95\%$ $50.88\%$ $68.65\%$ $63.28\%$ $94.63\%$
Table 6: LM baseline accuracy when the question is presented before vs after the items. Reported for SelectOne task for object categorization. All LMs were evaluated on the same draw of 1024 examples varying the question placement.
Models Ques Placement
Before After
Llama-70B $\% 99.3299.32\%$
Gemma-27B $9.41\% 99$$9.02\%$
Gemma-9B $8.93\% 99$$9.61\%$
Llama-8B $99.51\%$ $98.44\%$
Table 7: LM baseline accuracy for the SelectOne tasks that require dealing with different semantic information. All the LMs evaluated on the same draw of 1024 examples per semantic type. The question was presented after the options here.
Models Semantic Types
Object Profession Nationality Landmark-country
Llama-70B $99.32\%$ $8.44\% 99$$9.22\%$ $100.00\%$
Gemma-27B $99.02\%$ $98.93\%$ $91.60\%$ $98.14\%$
Gemma-9B $98.93\%$ $98.73\%$ $96.78\%$ $98.14\%$
Llama-8B $98.44\%$ $98.14\%$ $94.04\%$ $98.24\%$
@@ Appendix B Dual Implementation of Filtering in LMs: Question Before vs After
Section˜3.1c shows that filter heads are minimally causal when patched from question-before to question-before prompt, even when both follow the same prompt template and item presentation style. To understand this better, we perform a multi-step causal mediation analysis.
@@ 3. O3:
To test this hypothesis, we ablate the is_match flag in an item by replacing all of the residual stream latents of the tokens of that item with their corresponding latents cached for the same item in a neutral prompt (see Figure˜9 for details).
@@
When we perform this ablation for each of the items in the collection, we indeed see a significant drop in LM performance for the question-before setting, while the performance remains mostly unchanged for the question-after setting (see Table˜8).
This experiment supports our hypothesis that the question-before setting allows the LM to eagerly evaluate whether an item satisfies the predicate or not, store this intermediate result in the residual latents of the items as it processes the collection, and relies on that to make the final decision. We also notice that the question-after setting shows minimal sensitivity to this flag-ablation, which suggests that the processing of items do not rely on the context here: the LM populates the semantics (enrichment in Geva et al. (2023)) of each item in a context independent manner first, and then applies the predicate to perform filtering when the question is presented after.
@@
Our most decisive evidence comes from swapping the is_match flag between items: if we swap the is_match flag stored in $c_{\text{pos}}$ that satisfies the predicate with $c_{\text{neg}}$ that does not satisfy the predicate, we should expect the LM to change its answer from $c_{\text{pos}}$ to $c_{\text{neg}}$ if it relies on the is_match flag. To test this hypothesis, we set up another activation patching experiment (illustrated in Figure˜10, which is a more elaborate version of Figure˜5 in the main text). We perform a 2 part intervention:
@@ 2. I2:
Figure 11: Effect of swapping the is_match flag between items. The figures on the left shows the effect of patching only the residual states (Figure˜7) and figures on the right are when we additionally swap the is_match with another item (Figure˜10). The pair of figures on the top shows both cases in the question-before format (a). We observe that $c_{\text{flag}}$ becomes the top scoring item when the patching is performed in early layers. But this swapping of flags has no effect in the question after case, pair of figures on the bottom (b).
@@
We do not claim that the LM only relies on the is_match flag in the question-before setting. The fact that we see a bump in the score of $c_{\mathrm{targ}}$ ( blue) in Figure˜8 and patching $q_{\mathrm{src}}$ from question-before to question-after prompts has non-trivial causality (see Section˜3.1c) indicates that the LM does carry the predicate information in middle layers even in the question-before setting, although not as strongly as in the question-after setting.
This dual filtering implementation strategy — lazy evaluation via filter heads versus storing eager evaluation information with is_match flag — exemplifies a broader principle in neural computation: transformer LMs can maintain multiple pathways for core operations, dynamically selecting strategies based on what information is available. And the fact that filter heads still remain partially active even in the question-first setting shows that these mechanisms operate in parallel rather than in mutual exclusion.
@@ Appendix C Different Approaches for Locating the Filter Heads
$\mathrm{FilterScore}([\ell,j],\mathcal{C},\psi)$ $=\mathrm{score}_{\ell j}(c\mid\psi(c))-\max_{\neg\psi(c)}\;\big(\mathrm{score}_{\ell j}(c)\big)$ (6)
$\text{where,}\quad\mathrm{score}_{\ell j}(c)$ $=\sum_{t\in c}\mathrm{Attn}_{[\ell,j]}(q_{-1},K)_{t}$
@@
Figure 12: Collective causality of 79 filter heads identified with FilterScore, CMA, and CMA with learned masks. Evaluated on the same 512 examples from the SelectOne task
@@
Activation Patching.
We can patch the attention heads individually and quantify their indirect effect at mediating the target property in the patched run. Todd et al. (2024) and Feucht et al. (2025) identified Function Vector heads and Concept heads with this approach. Specifically, for each head $\ell,j]$ we patch $q_{\mathrm{src}}$ from the source run to the destination run with the method detailed in Section˜2 and check its effect on boosting the score (logit) of the target item, $c_{\mathrm{targ}}$. The indirect effect is measured as $\mathrm{logit}[\leftarrow q_{\mathrm{src}}]\big(c_{\mathrm{targ}}\big)-\mathrm{logit}\big(c_{\mathrm{targ}}\big)$.
Activation Patching with Learned Masks.
In our analysis we find that patching a single head is not enough to causally influence the LLM behavior. This is because the same mechanism can have parallel implementations (McGrath et al., 2023), which may work against the intervention to rectify its effects. To overcome this issue, we learn a differential mask to locate the minimal set of components (attention heads) that collectively implement filtering. Following De Cao et al. (2020); Davies et al. (2023); and Csordás et al. (2021), we optimize a differential binary mask over all of the attention heads. The idea is to learn a binary mask for the neural components (query states of attention heads in our case) such that after an interchange intervention (Eq.˜3) is performed, the LM shows the desired counterfactual behavior (executing the predicate from the source prompt).
When we compare different head identification approaches while controlling for the number of heads selected, the learned mask approach achieves the highest causality (Figure˜12).
Figure 13: Location of filter heads in Llama-70B. (a) shows the individual FilterScore for each head: how much they attend to the correct option over others. (b) shows the indirect effect: how much patching $q_{\mathrm{src}}$ from a single head promotes the predicate target. The filter heads identified with Section˜2.3 are marked with black borders.